OpenStudy (australopithecus):

I have the problem x = e^(-t)cos(2012t) y = e^(-t)(2012t) 0≤ t ≥pi calculate its exact length. so I know I have to use the therom

OpenStudy (australopithecus):

\[\int\limits_{a}^{b}\sqrt{(dx/dt)^{2} + (dx/dt)^{2}}dx\]

OpenStudy (australopithecus):

I have so far that dx/dt = -e^(-t)cos(2012t) - e^(-t)sin(2012t)2012 and dy/dt = -e^(-t)sin(2012t)+e^(t)cos(2012t)2012

OpenStudy (australopithecus):

the integral I'm getting is very very ugly I simplified it to \[\int\limits_{0}^{pi}\sqrt{\sin^{2}(2012t)2012^{2} + \cos^{2}(2012t) -\sin^{2}(2012t) + \cos^{2}(2012t)2012^{2}}dx/e^{t}\]

OpenStudy (australopithecus):

still very ugly :l should I have used quotient rule instead?

OpenStudy (australopithecus):

I simplified it down to \[\int\limits_{0}^{pi} \sqrt{2012^{2} - 1 + 2\cos^{2}(2012t)}dx/e^{t}\]

OpenStudy (dumbcow):

shouldn't the dx be dt

OpenStudy (australopithecus):

or \[\int\limits_{0}^{pi} \sqrt{4048143 - 2\cos{2}(2012t)}dt/e^{t}\] good call dumbcow :)

OpenStudy (australopithecus):

It says it should be easy after a simplification :\ i must have did something wrong or he lied

OpenStudy (dumbcow):

ok i think everything checks out until last step what happened to sin^2 ?

OpenStudy (australopithecus):

cos^2(2012t)−sin^2(2012t) = 2cos^(2)(2012t)

OpenStudy (australopithecus):

by the trig rule sin^(2)(x) + cos^(2)(x) = 1

OpenStudy (australopithecus):

maybe that simplification made it less simplified :l

OpenStudy (dumbcow):

oh right i didn't see the "-1" then i was thinking of identity: cos(2x) = cos^2 - sin^2

OpenStudy (australopithecus):

its funny because my prof always said he was evil now I know he wasn't lying

OpenStudy (australopithecus):

such a long and tedious integral it seems although I may just be dumb

OpenStudy (australopithecus):

*might

OpenStudy (dumbcow):

wait, tell me how you factored the e^t out of sqrt sorry im doing this in my head right now

OpenStudy (australopithecus):

yes (e^(-2t))^(1/2) = (1/2)^(1)/(e^(2t))^(1/2) = 1/2^(t)

OpenStudy (australopithecus):

sorry that should be 1/e^(t)

OpenStudy (dumbcow):

no isn't attached to a sin*cos when you square dx/dt and dy/dt

OpenStudy (dumbcow):

wow nevermind

OpenStudy (australopithecus):

?

OpenStudy (australopithecus):

I could try u substitution for this but it will still be really dirty

OpenStudy (australopithecus):

due to the e^(t)

OpenStudy (australopithecus):

maybe integration by parts but even then that will probably even be grosser

OpenStudy (dumbcow):

no i wouldn't recommend integration by parts

OpenStudy (dumbcow):

lets see, its in the form \[\int\limits_{ }^{} \frac{\sqrt{u^{2}+\cos(2u)}}{e^{t}} dt\] where u = 2012 right

OpenStudy (australopithecus):

that actually looks a lot better lol

OpenStudy (dumbcow):

i have no idea what to do...haha lets see what wolfram can do http://www.wolframalpha.com/input/?i=integrate+sqrt%281%2Bcos%282t%29%29%2Fe^t+dt

OpenStudy (australopithecus):

then we have \[\int\limits_{0}^{pi} \sqrt{u^{2} + \cos^{2}(u) - \sin^{2}(u)-1}du/e^{t}\]

OpenStudy (australopithecus):

you forgot to include -1

OpenStudy (australopithecus):

or + 1 rather

OpenStudy (australopithecus):

no it was -1

OpenStudy (dumbcow):

hmm cos(2x) = cos^2 - sin^2 right

OpenStudy (australopithecus):

right

OpenStudy (australopithecus):

man this integral is frustrating i need to eat something brb

OpenStudy (dumbcow):

from first sqrt \[\rightarrow \sqrt{u^{2}(\sin^{2} + \cos^{2}) + \cos^{2} -\sin^{2}} = \sqrt{u^{2} + \cos(2ut)}\] the "-1" only comes into play if you use other form cos^2 - sin^2 = 2cos^2 - 1

OpenStudy (australopithecus):

so i made a mistake?

OpenStudy (australopithecus):

It has to be integrable

OpenStudy (dumbcow):

here is definite integral ...sorry i meant it has no known anti-derivative http://www.wolframalpha.com/input/?i=integrate+sqrt%282012^2%2Bcos%282%282012%29t%29%29%2Fe^t+dt+from+0+to+pi

OpenStudy (australopithecus):

you must have made a mistake

OpenStudy (dumbcow):

why do you have x's in there ?

OpenStudy (australopithecus):

it can be solved

OpenStudy (dumbcow):

ahh i went back with paper and checked everything, you got a sign mixed up somewhere notice on wolfram answer...all those (sin^2 +cos^2) can be replaced with 1 integral can be reduced to \[\sqrt{2012^{2} +1}\int\limits_{0}^{\pi} e^{-t} dt\]

OpenStudy (dumbcow):

the cos(2012)^2 - sin(2012)^2 should be cos(2012)^2 + sin(2012)^2

OpenStudy (australopithecus):

when I was expanding?

OpenStudy (australopithecus):

oh nvm I think I see thanks :)

OpenStudy (dumbcow):

whats funny is you get the same answer of 1925.05 either way interesting problem your professor came up with

OpenStudy (anonymous):

a=2012 \[x(t) =e^{-t} \cos (a\,t)\\ y(t)= e^{-t} \sin (a\, t)\\ \] Is this our curve?

OpenStudy (anonymous):

\[ x'(t)^2+y'(t)^2 =\left(a e^{-t} \cos (a t)-e^{-t} \sin (a t)\right)^2+\left(-a e^{-t} \sin (a t)-e^{-t} \cos (a t)\right)^2=\\ \left(a^2+1\right) e^{-2 t}\\ \sqrt{ x'(t)^2+y'(t)^2 }= \sqrt{ 1 + a^2} e^{-t} \]

OpenStudy (anonymous):

\[ \sqrt{1 + a^2} \int_0^\pi e^{-t} dt=\sqrt{1 + a^2}(1 - e^{-\pi})\approx 1925.05 \]